7 + X/2 = 10
X/3 - 5 = -7
Please explain how you worked it on at least one of them. Thank you

Answer by rfer(16322)

7+x/2=10
x/2=10-7
x=3*2
x=6
----------
x/3-5=-7
x/3=-7+5
x=-2*3
x=-6